# Learning Object State Changes in Videos: An Open-World Perspective

### v1

@@ 5 Experiments
Datasets In addition to our new HowToChange dataset, we also evaluate on ChangeIt [49] due to its expansive data scale (34K videos spanning many activities) and high relevance to our task. Beyond the conventional split of ChangeIt, we propose a new split tailored to our open-world formulation. Specifically, from the 44 available OSC categories in ChangeIt, we concentrate on the 25 categories where each state transition is paired with more than one object. With those, we form the “ChangeIt (open-world)” subset that comprises 8 state transitions and 25 objects. Within each state transition, objects are randomly divided into known and novel categories, yielding 13 known OSC categories for training and 12 novel OSC categories exclusively reserved for evaluation. A detailed breakdown of this split can be found in Supplementary. To sum up, our evaluation encompasses: ChangeIt; ChangeIt (open-world); and HowToChange, offering a comprehensive setup in both closed-world and open-world scenarios.
@@
Implementation Videos are sampled at one frame per second. Each one-second video segment gets assigned to an OSC state label, and is encoded by InternVideo [58], a general video foundation model (which is kept frozen for training efficiency). Our video model consists of a 3-layer transformer with a hidden dimension of 512 and 4 attention heads as the encoder and a 1-layer MLP as the decoder. Consistent with prior works [49, 50], our model is designed to predict the OSC state label (i.e., initial, transitioning, end state, or background) within a given video. While the standard output does not include state transition name, our multi-task version detailed in Supp. is capable of this. See Supp. for further discussion on this and complete evaluation, baseline and implementation details.
Main Results Table 2 presents results on the three datasets. VidOSC outperforms all approaches in both traditional closed-world and our newly introduced open-world settings. The large performance gains underscore the effectiveness of two pivotal components within VidOSC. First, its use of text and VLMs for supervisory signals during training greatly boosts performance. Particularly on novel OSCs not encountered during training, VidOSC extracts meaningful cues from the video modality to refine pseudo labels from VLMs, and ultimately surpass the VLM baselines. Second, our model for object-agnostic state prediction, designed for the open-world context, effectively narrows the gap between known and novel OSCs. For instance, on ChangeIt (open-world), MultiTaskChange [50] experiences a 19% decline in state precision@1 while VidOSC has only a 8% drop. Note that we observe a more pronounced performance drop of all approaches on ChangeIt (open-world) than on HowToChange due to its limited number of objects available per state transition. These results also point to potential for future development in bridging the known and novel OSC performance gap.
Qualitative Results Fig. 4 presents VidOSC’s top-1 frame predictions on ChangeIt (open-world) and HowToChange, across both known and novel OSCs. Notably, despite never seeing objects such as cauliflower and capsicum during training, VidOSC effectively leverages visual state change cues and correctly localizes the three states of these objects going through OSCs, demonstrating its strong generalization capability. Transitioning to a more holistic view, Fig. 5 compares VidOSC’s frame-by-frame predictions with all baselines for a given test video. VidOSC provides temporally coherent predictions, smoothly progressing through the OSC states in the natural order (i.e., from initial to transitioning then end). In contrast, baseline approaches often yield fragmented and inconsistent predictions, indicating a lack of understanding of the video’s global temporal context, primarily due to their reliance on frame-wise modeling. See Supp. and Supp. video for more qualitative examples and VidOSC’s interpretability on object relations.
@@
Ablation To further dissect the performance gains brought by our three model design techniques (i.e., shared state vocabulary, temporal modeling and object-centric features), we conduct an ablation study, removing one component at a time. The results, presented in Table 3, confirm the essential role of each element. A shared state vocabulary is particularly crucial in the open-world context, as its absence increases the gap between known and novel OSCs from 2.5% to 5.2%. Furthermore, temporal modeling provides a substantial performance boost, and object-centric features offer further gains. See Supp. for an additional analysis of VidOSC’s performance with different pseudo labels.
@@
Figure 6: Frame retrieval on the HowToChange test set. Given a query frame showcasing a novel OSC at its initial state, VidOSC retrieves the nearest and furthest frame among all known OSCs at their end states. The closet post-OSC frame follows the transition trajectory of the query while the furthest post-OSC frame depicts a substantially different object, demonstrating VidOSC’s capability to generalize the OSC progression for novel objects.
@@ 6 Conclusion
This work aims at a comprehensive exploration of video OSCs, with a novel open-world formulation. To address the challenges, we leverage text and VLMs to assist the training of a video OSC model at scale and design three modeling techniques to achieve object-agnostic state prediction for better generalization to novel OSCs. Furthermore, we present the most expansive video OSC dataset collection HowToChange, that echoes the natural long-tail of state transitions coupled with varied objects, fostering a realistic representation of real-world scenarios. As for future work, we consider extending VidOSC to video sequences featuring concurrent OSC processes, and integrating a spatial understanding of OSC within our open-world framework.

### v2

@@ 5 Experiments
Datasets In addition to our new HowToChange dataset, we also evaluate on ChangeIt [50] due to its expansive data scale (34K videos spanning many activities) and high relevance to our task. Beyond the conventional split of ChangeIt, we propose a new split tailored to our open-world formulation. Specifically, from the 44 available OSC categories in ChangeIt, we concentrate on the 25 categories where each state transition is paired with more than one object. With those, we form the “ChangeIt (open-world)" subset that comprises 8 state transitions and 25 objects. Within each state transition, objects are randomly divided into known and novel categories, yielding 13 known OSC categories for training and 12 novel OSC categories exclusively reserved for evaluation. A detailed breakdown of this split can be found in Supp. To sum up, our evaluation encompasses: ChangeIt; ChangeIt (open-world); and HowToChange, offering a comprehensive setup in both closed-world and open-world scenarios.
@@
Implementation Videos are sampled at one frame per second. Each one-second video segment gets assigned to an OSC state label, and is encoded by InternVideo [59], a general video foundation model (which is kept frozen for training efficiency). Our video model consists of a 3-layer transformer with a hidden dimension of 512 and 4 attention heads as the encoder and a 1-layer MLP as the decoder. Consistent with prior work [50, 51], our model predicts the OSC state label (i.e., initial, transitioning, end state, or background) for a video, assuming the video OSC category is known (say from a recognition model’s output or a user’s specific query). While the standard output does not include the state transition name, our multi-task version detailed in Supp. is capable of this.
Main Results Table 2 presents results on all three datasets. VidOSC outperforms all approaches in both traditional closed-world and our newly introduced open-world settings. The large performance gains—as much as 9.6% jumps in precision vs. the next best method—underscore the effectiveness of two pivotal components in VidOSC. First, its use of text and VLMs for supervisory signals during training: particularly on novel OSCs , VidOSC extracts meaningful cues from the video modality to refine pseudo labels from VLMs, and ultimately surpasses the VLM baselines. Second, our model for object-agnostic state prediction, designed for the open-world context, effectively narrows the gap between known and novel OSCs. For instance, on ChangeIt (open-world), MultiTaskChange [51] experiences a 19% decline in state precision@1 while VidOSC has only a 8% drop. Note that we observe a more pronounced performance drop of all approaches on ChangeIt (open-world) than on HowToChange due to its limited number of objects available per state transition. These results also point to the potential for future development in bridging the known and novel OSC performance gap.
Qualitative Results Fig. 4 presents VidOSC’s top-1 frame predictions on ChangeIt (open-world) and HowToChange, across both known and novel OSCs. Notably, despite never seeing objects such as cauliflower and capsicum during training, VidOSC effectively leverages visual state change cues and correctly localizes the three states of these objects going through OSCs, demonstrating its strong generalization capability. For a more holistic view, Fig. 5 compares VidOSC’s frame-by-frame predictions with all baselines for a given test video. VidOSC provides temporally coherent predictions, smoothly progressing through the OSC states in the natural order (i.e., from initial to transitioning then end). In contrast, baseline approaches often yield fragmented and inconsistent predictions, indicating a lack of understanding of the video’s global temporal context, primarily due to their reliance on frame-wise modeling. See Supp. and Supp. video for more qualitative examples and VidOSC’s interpretability on object relations.
@@
Ablation To further dissect the performance gains brought by our three model design techniques (i.e., shared state vocabulary, temporal modeling, and object-centric features), we conduct an ablation study, removing one component at a time. Table 3 confirms the essential role of each element. A shared state vocabulary is particularly crucial in the open-world context, as its absence increases the gap between known and novel OSCs from 2.5% to 5.2%. Furthermore, temporal modeling provides a substantial performance boost, and object-centric features offer further gains. See Supp. for an additional analysis of VidOSC’s performance with different pseudo labels.
@@
Figure 6: Frame retrieval on the HowToChange test set. Given a query frame showcasing a novel OSC at its initial state, VidOSC retrieves the nearest and furthest frame among all known OSCs at their end states. The closest post-OSC frame follows the transition trajectory of the query while the furthest post-OSC frame depicts a substantially different object, demonstrating VidOSC’s capability to generalize the OSC progression for novel objects.
@@ 6 Conclusion
This work aims at a comprehensive exploration of video OSCs, with a novel open-world formulation. To address the challenges, we leverage text and VLMs to assist the training of a video OSC model at scale and design three modeling techniques to achieve object-agnostic state prediction for better generalization to novel OSCs. Furthermore, we present the most expansive video OSC dataset collection HowToChange, which echoes the natural long-tail of state transitions coupled with varied objects, fostering a realistic representation of real-world scenarios. As for future work, we will consider extending VidOSC to video sequences featuring concurrent OSC processes, and integrating spatial understanding of OSC within our open-world framework.
Acknowledgements: UT Austin is supported in part by the IFML NSF AI Institute. KG is paid as a research scientist by Meta.